OpenStudy (anonymous):

Determine the type of boundary line and shading for the graph of the inequality 4x + y > -6. Dashed line with shading on the side that includes the origin. Solid line with shading on the side that does not include the origin. Dashed line with shading on the side that does not include the origin. Solid line with shading on the side that includes the origin.

OpenStudy (cruffo):

Solve for y: y > -4x - 6 When ever you have: > or < use a dashed line \(\leq\) or \(\geq\) use a solid line When you have y > or \(y \geq\) shade above y< or \(y\leq\) shade below

OpenStudy (anonymous):

ok thank u!! would this be correct to solve for y? y > -4x - 6 + 4 +4 y > -2

OpenStudy (cruffo):

no... when you have a bunch of letter, you won't get the typical y = some number. In this case, "solve for y" is done when you get y by itself. So y > -4x-6 is "solved for y".

OpenStudy (anonymous):

oh omg i forgot all about the x

OpenStudy (cruffo):

lol